### v1

# AWM[Uncaptioned image]: Accurate Weight-Matrix Fingerprint for Large Language Models

Abstract
Protecting the intellectual property of large language models (LLMs) is crucial, given the substantial resources required for their training. Consequently, there is an urgent need for both model owners and third parties to determine whether a suspect LLM is trained from scratch or derived from an existing base model. However, the intensive post-training processes that models typically undergo—such as supervised fine-tuning, extensive continued pretraining, reinforcement learning, multi-modal extension, pruning, and upcycling—pose significant challenges to reliable identification. In this work, we propose a training-free fingerprinting method based on weight matrices. We leverage the Linear Assignment Problem (LAP) and an unbiased Centered Kernel Alignment (CKA) similarity to neutralize the effects of parameter manipulations, yielding a highly robust and high-fidelity similarity metric. On a comprehensive testbed of 60 positive and 90 negative model pairs, our method demonstrates exceptional robustness against all six aforementioned post-training categories while exhibiting a near-zero risk of false positives. By achieving perfect scores on all classification metrics, our approach establishes a strong basis for reliable model lineage verification. Moreover, the entire computation completes within 30s on an NVIDIA 3090 GPU.
1 Introduction
Figure 1: On 150 LLM pairs, our method achieves perfect AUC in distinguishing related and unrelated LLMs, outperforming baselines.
Large language models (LLMs) have become foundational to many artificial intelligence applications. However, training an LLM from scratch demands substantial computational resources and vast amounts of data. Consequently, most open-source models are released under specific licenses (Touvron et al., 2023a; Team et al., 2025; Mesnard et al., 2024; Kamath et al., 2025) or require an application (Touvron et al., 2023b; Zhang et al., 2022; Penedo et al., 2023; BaiChuan-Inc, 2023; Team, 2023; Zheng et al., 2023; Grattafiori et al., 2024) and approval process to protect intellectual property. Despite these measures, some developers may circumvent such protections by wrapping or post-training existing base LLMs, then falsely claiming to have trained their own models. Recent controversies (pzc163 et al., 2024; Yoon et al., 2025) have underscored the urgent need to determine whether a suspect model is genuinely trained from scratch or derived from an existing base model.
The core challenge lies in extracting a stable fingerprint to identify the true base model. This task is complicated by the fact that LLMs often undergo heavy post-training processes, such as supervised fine-tuning (SFT), extensive continued pretraining (Team et al., 2024; Hui et al., 2024), and reinforcement learning (RL). Furthermore, emerging techniques like extension to multimodal tasks, architectural pruning (Xia et al., ), and upcycling (He et al., 2024) can drastically alter a model’s parameters, outputs, and even its structure, posing substantial challenges for identification. Moreover, malicious actors might intentionally manipulate weight matrices through operations including scaling, permutation, pruning, and even rotation to obscure a model’s origin.
Therefore, a practical fingerprinting method needs to satisfy the following critical conditions:
  • Robustness against extensive post-training processes.
  • Resilience to malicious weight manipulations such as scaling, pruning, permutation, and rotation.
  • Performance Preservation, ensuring no degradation of the LLM’s capabilities, as both users and manufacturers place a high premium on model performance.
  • High Fidelity, possessing sufficient discriminative power and an extremely low false-positive rate to prevent false accusations of model theft.
  • Computational Efficiency, remaining lightweight enough for comparisons given the immense parameter counts of modern LLMs.
A review of previous work reveals that existing methods often fail to meet one or more of these criteria. Watermarking techniques, for instance, embed identifiable signals via extra training (Peng et al., 2023; Xu et al., 2024), but this can degrade performance (Russinovich & Salem, 2024), may not survive aggressive post-training (Fernandez et al., 2024; Gubri et al., 2024), and must be applied before the model’s release. Other existing fingerprinting methods either suffer from high false-positive rates (Zhang et al., 2025) or lack robustness against heavy post-training modifications like continued pretraining (Zeng et al., 2024), a limitation we confirm in our experiments.
In this work, we propose a novel approach that satisfies all the aforementioned requirements. Our method begins with an analysis of LLM weight manipulations. By leveraging the Linear Assignment Problem (LAP) and an unbiased Centered Kernel Alignment (CKA), we derive a similarity metric that is robust to all these manipulations. Our entire similarity computation completes within 30 seconds on a single NVIDIA 3090 GPU. Furthermore, our method is training-free and does not impair the LLM’s performance.
We validate our approach on a comprehensive test set of 60 positive (base-offspring) and 90 negative (independent) model pairs. Compared to state-of-the-art methods such as HuRef and REEF, our approach achieves a much larger separation gap while maintaining a near-zero false-positive risk. Crucially, our method is the only one to prove robust against all tested forms of post-training: SFT, extensive continued pretraining (up to 5.5T tokens), reinforcement learning, multi-modal extension, pruning, and upcycling. On this 150-pair dataset, our method achieves a perfect Area Under the Curve (AUC), partial AUC (for False Positive Rate <0.05<0.05), and True Positive Rate @ 1% False Positive Rate of 1.0, establishing a strong basis for reliable and robust model lineage verification.
2 Related Work
Model copyright protection methods fall into two main categories: watermarking and fingerprinting.
Watermarking. Watermarking methods typically involve finetuning models to inject backdoor triggers that prompt the model to generate predefined content, or embedding watermarks directly into model weights for identification purposes. A substantial body of research has explored watermarking for smaller DNNs such as CNNs and BERT (Devlin et al., 2019), including encoding watermarks into model weights (Chen et al., 2019a; Wang & Kerschbaum, 2021; Liu et al., 2021; Uchida et al., 2017) and injecting triggers to produce specific outputs (Adi et al., 2018; Guo & Potkonjak, 2018; Le Merrer et al., 2020; Chen et al., 2019b). However, these methods are often task-specific and not well-suited for foundation LLMs. For watermarking LLMs, researchers have proposed various methods to inject watermarks for identification (Russinovich & Salem, 2024; Xu et al., 2024; Li et al., 2023; Peng et al., 2023; Kirchenbauer et al., 2023; Zhao et al., 2023). Nevertheless, these approaches inevitably compromise LLM performance and are not robust to extensive post-training.
Fingerprinting. Fingerprinting methods extract intrinsic model features as fingerprints for identification, without requiring additional training and thus without impacting model performance. For small DNNs, prior works (Zhao et al., 2020; Pan et al., 2022; Yang et al., 2022; Lukas et al., 2019; Peng et al., 2022) analyze model behaviors on preset test cases to construct fingerprints. For LLMs, HuRef (Zeng et al., 2024) is a classic weight-based fingerprinting method that derives invariant terms and computes similarities between them for model identification. However, it is not robust to extensive training. REEF (Zhang et al., 2025), from the perspective of representation, measures similarities between model representations, but suffers from a high false positive rate, which limits its practical use. To date, there is still a lack of fingerprinting methods for LLMs that are both robust to extensive training and exhibit a very low risk of false positives.
3 Preliminary
LLM Architecture
Most Large Language Models (LLMs) follows a decoder-only Transformer architecture (Radford et al., 2018). The details of various LLMs may differ, yet the Transformer blocks are similar. In particular, a Transformer block in an LLM usually consists of residual connections (He et al., 2016), Root Mean Square Normalization (RMSNorm, Zhang & Sennrich (2019)), the self-attention mechanism (Lin et al., 2017), Rotary Position Embedding (RoPE, Su et al. (2024)), and a feed-forward network (FFN). We denote by $\mathbb{W}_{A}$ the set of an $L$-layered LLM A’s weights, and $\mathbb{W}_{A,\text{partial}}=\{W_{A,\text{emb}}\}\bigcup_{l=1}^{L}\{W_{A,i}^{(l)}\mid i\in\{\text{Q},\text{K}\}\}$ the set of word embeddings and $\text{Q},\text{K}$ matrices, where $W_{A,\text{emb}}$ is the word embeddings and $W_{A,\text{Q}}^{(l)},W_{A,\text{K}}^{(l)}$ are the $\text{Q},\text{K}$ matrices at the $l$-th layer. We defer the rest of the notations to Appendix A.1.
Central Kernel Alignment (CKA)
Proposed in Kornblith et al. (2019), CKA is a similarity detection method (Zhang et al., 2025) based on Hilbert-Schmidt Independence Criterion (HSIC, Gretton et al. (2005)). It is invariant to column-wise orthogonal transformations and constant multiplications:
Theorem 3.1.
For any input matrices $X_{1}\in\mathbb{R}^{m\times n_{1}},X_{2}\in\mathbb{R}^{m\times n_{2}}$, any orthogonal transformations $U_{1}\in\mathbb{R}^{n_{1}\times n_{1}},U_{2}\in\mathbb{R}^{n_{2}\times n_{2}}$, any non-zero constants $c_{1},c_{2}\in\mathbb{R}$,
$\text{CKA}(X_{1},X_{2})=\text{CKA}(c_{1}X_{1}U_{1},c_{2}X_{2}U_{2}).$ (1)
Even with semi-orthogonal $U_{1}$ and $U_{2}$ where $U_{1}\in\mathbb{R}^{n_{1}\times n_{1}^{\prime}}$, $U_{2}\in\mathbb{R}^{n_{2}\times n_{2}^{\prime}}$, $n_{1}>n_{1}^{\prime}$, $n_{2}>n_{2}^{\prime}$, CKA still preserves input similarity to some extent (Kang et al., 2025; Chun et al., 2025). Nevertheless, the standard HSIC estimator converges at rate $1/\sqrt{m}$ and shows finite-sample bias (Gretton et al., 2005; Murphy et al., 2024). To this end, an unbiased version (Song et al., 2007) is proposed and extends the value of CKA from $[0,1]$ to [1,1][-1,1]. Additionally, linear kernels are often selected in CKA due to their computational efficiency and similar performance to other kernels (Kornblith et al., 2019). We give the formal definition of CKA in Appendix A.2, and the proof of Theorem 3.1 in Appendix B.1.
4 Manipulations on An Open-Source LLM
4.1 Problem Definition
Weights of an open-source LLM are often inherited, but unclaimed inheritance invites manipulations. Although the source code may not explicitly reveal these manipulations, the weights are vulnerable to modifications that leave no trace in code (e.g., post-hoc matrix multiplications to produce new weights), and they can even be altered to evade independence tests. To enable detection on model weights, we first formalize the plausible manipulation forms.
Definition 4.1 (LLM Weight Manipulations).
Let A and B be two open-source LLMs with the same number of layers $L$, and $\mathbb{W}_{A},\mathbb{W}_{B}$ denote the sets of matrix weights for A and B respectively. Then, if B manipulates A, the manipulations on $W_{A,\text{partial}}$ are
$W^{(l)}_{B,i}=L^{(l)}_{B,i}W^{(l)}_{A,i}R^{(l)}_{B,i}+E^{(l)}_{B,i}\quad\text{for all }1\leq l\leq L\text{ and }i\in\left\{\text{Q},\text{K}\right\}$ (2)
$W_{B,\text{emb}}=W_{A,\text{emb}}R_{B,\text{emb}}+E_{B,\text{emb}}$ (3)
where $L$ and $R$ are row-wise and column-wise transformation matrices, and $E$ is the error term. The learnable weights in RMSNorm are also changed correspondingly.
We omit the row-wise transform on word embeddings, $L_{B,\text{emb}}$, since each row of $W_{A,\text{emb}}$ represents a token and mixed token representations are hard to be faithfully recovered in the calculation of attention scores. We further assume that the suspicious and manipulated models produce similar (or identical) outputs to be consistent with the goal of reusing base model performance. This assumption induces constraints on the transformation matrices and, in turn, enables detection via weight-matrix similarity tests. We first develop a fine-grained view of the constraints on $W_{A,\text{partial}}$ in what follows.
4.2 Residual Connections: Passing Manipulations Forward
A Transformer block consists of multiple components linked by residual connections. If one component takes input $X$, the next receives $Y=X+f(X)$, so any manipulation $T$ on $X$ must be recovered within the component to propagate as the next input (Zeng et al., 2024). This propagation also depends on the constituent functions of the component, which either commute with the manipulation to recover it or remain invariant and absorb it. We formalize this in the following:
Proposition 4.2 (Proof in Appendix B.2).
Let $f=f_{n}\circ\dots\circ f_{1}$ be a component in Transformer and let $T$ be a manipulation in the input. If for every $k\in\{1,\dots,n\}$, $T\circ f_{k}=f_{k}\circ T$, then $f\circ T=T\circ f$, i.e. the manipulation propagates through constituent functions of a component.
Compared to Zeng et al. (2024), Section 4.2 shows that residual connections are more vulnerable under a component-wise view, since constituent functions can propagate manipulations in various ways. However, nonlinearities of the functions, especially those within the self-attention mechanism, pose constraints for manipulations on both inputs and weights. We next split the self-attention mechanism into two constituent functions, RMSNorm and attention (see Section A.1), and show how these effects arise and are constrained these two functions with a focus on RMSNorm, RoPE and attention scores.
4.3 RMSNorm: A Constraint on Embedding Manipulations
Any input to the self-attention mechanism, including word embeddings and hidden states, first go through RMSNorm. However, RMSNorm can facilitate potential manipulations on inputs, because it commutes with certain transformations $R_{B,\text{emb}}$ of the embeddings:
Theorem 4.3.
Let models $A$ and $B$ share the same [^1] The conclusions generalize to models with different number of layers (possibly a result of pruning). See Appendix B.3architecture. Let $c\neq 0$ be a scalar, $P$ be a (partial) permutation matrix, and $D$ be a signature matrix. Then, $R_{B,\text{emb}}=cPD$ can be recovered after RMSNorm in model B if related RMSNorm parameters are adjusted.
Theorem 4.3 indicates that RMSNorm is susceptible to embedding manipulations composed of constant multiplications, permutations and sign flips. On the other hand, other manipulations on word embeddings are generally neither commutative with RMSNorm nor invariant to it, which potentially brings a constraint to manipulations. We provide a proof for Theorem 4.3 in Appendix B.3, along with a discussion on other manipulations on word embeddings.
4.4 RoPE and Attention Scores: Boundaries for Q/K Manipulations
After RMSNorm, inputs are fed into attention score calculations. The nonlinearity here, particularly in the softmax and RoPE functions, poses a barrier for manipulations on inputs to pass through. Hence, we follow Zeng et al. (2024) to assume that input manipulations does not change the value of attention scores. The manipulations on Q,K matrices are thus constrained under this assumption.
Theorem 4.4.
The manipulations on Q,K matrices at layer $l$ can be categorized into
  1. 1.
    Input-related ones, passed by RMSNorm: $W_{B,i}^{(l)}=c^{-1}W_{A,i}^{(l)}PD$,  for $i\in\{\text{Q,K}\}$;
  2. 2.
    RoPE-related ones: $W_{B,i}^{(l)}=U_{B,i}^{(l)}W_{A,i}^{(l)}$,  for $i\in\{\text{Q,K}\}$;
where $U_{B,i}^{(l)}$ are special orthogonal matrices that keep RoPE results.
A proof for Theorem 4.4 is provided in Appendix B.4, where we also show how manipulations on inputs are recovered after V, O matrices to satisfy Section 4.2. These manipulations preserve the attention score values of model A in the suspicious model B. However, they can greatly change the weights of the original model A, bringing difficulty to the development of detection methods.
4.5 Potential Attacks
Combining Theorem 4.3 and Theorem 4.4, the admissible manipulations on $W_{A,\text{partial}}$ (word embeddings and $\text{Q},\text{K}$ matrices) in Section 4.1 are restricted to
$W_{B,i}^{(l)\top}=c^{-1}\,D^{\top}P^{\top}\,W_{A,i}^{(l)\top}\,U_{B,i}^{(l)\top}+E_{B,i}^{(l)\top},\quad 1\leq l\leq L,\;\;i\in\{\text{Q},\text{K}\},$ (4)
$W_{B,\text{emb}}=c\,W_{A,\text{emb}}\,PD+E_{B,\text{emb}}.$ (5)
Here $E$ collects post-training, continual pre-training, pruning, upcycling, multimodal adaptation, or related adjustments. The nonlinear usage of $W_{A,\text{partial}}$, especially through the $\text{Q},\text{K}$ matrices, substantially limits manipulation complexity. Consequently, checking similarity between $W_{A,\text{partial}}$ and $W_{B,\text{partial}}$ is typically sufficient for detection. The remaining avenues targeting $\mathbb{W_{A}}/\mathbb{W}_{A,\text{partial}}$ are deferred to Appendix B.5, where we also discuss how such transformations can recover Transformer block outputs in light of Section 4.2.
5 Methodology
Building on Theorem 4.3 and Theorem 4.4, we propose a two-stage procedure in 1: (i) extract $P$ and $D$ from the word embeddings; (ii) assess cross-model similarity by comparing the Q and K matrices. This design achieves fast, reliable discrimination with minimal computation.
Algorithm 1 LAP-Enhanced Unbiased Central Kernel Alignment (UCKA) Similarity Detection
1:Given two LL [^2] For LLMs with differing layer counts ($L_{A}$, $L_{B}$), e.g., from layer pruning, we find the optimal layer pairing by solving the LAP on an $L_{A}\times L_{B}$ matrix of layer-wise similarities before calculating overall similarity.-layered LLMs $A,B$ and their weight matrices $\mathbb{W}_{A,\text{partial}},\mathbb{W}_{B,\text{partial}}$.
2:Let $I=\text{Vocab}(A)\cap\text{Vocab}(B)$, and $m^{\prime}=|I|$.
3:Let $W_{A,\text{shared\text{-}emb}}=W_{A,\text{emb}}[I,:]$ and $W_{B,\text{shared\text{-}emb}}=W_{B,\text{emb}}[I,:]$.
4:Build the cosine similarity matrix $C$ with $C_{k,l}=\dfrac{\langle(W_{A,\text{shared\text{-}emb}})_{:,k},\,(W_{B,\text{shared\text{-}emb}})_{:,l}\rangle}{\|(W_{A,\text{shared\text{-}emb}})_{:,k}\|\,\|(W_{B,\text{shared\text{-}emb}})_{:,l}\|}$.
5:Construct the permutation and signature matrices $P,D$ via LAP:
6:  Find a permutation $\pi$ maximizing $\sum_{k}|C_{k,\pi(k)}|$ with the Hungarian algorithm.
7:  For each column $k$, set $s_{k}=\text{sign}\!\big(C_{k,\pi(k)}\big)$.
8:  Set $P_{k,\pi(k)}=1$ for every $k$, set $D=\text{diag}(s_{1},s_{2},\dots,s_{k},\dots)$.
9:for $l=1,\dots,L$ do
10:  $s_{Q}^{(l)}=\text{UCKA}\!\left(D^{\top}P^{\top}W^{(l)\top}_{A,Q},\,W^{(l)\top}_{B,Q}\right)$, $s_{K}^{(l)}=\text{UCKA}\!\left(D^{\top}P^{\top}W^{(l)\top}_{A,K},\,W^{(l)}_{B,K}\right)$
11:end for
12:return $\sum_{l=1}^{L}(|s_{Q}^{(l)}|+|s_{K}^{(l)}|)/2L$
Extracting the Permutation and Signature from Word Embeddings
A key property of $R_{B,\text{emb}}$ is that the permutation and signature matrices may appear in either order: both $R_{B,\text{emb}}=cPD$ and $R_{B,\text{emb}}=cDP$ are possible. However, any product $DP$ can be rewritten as $P^{\prime}D^{\prime}$ for some permutations $P^{\prime}$ and signature matrices $D^{\prime}$. Hence it suffices to recover a canonical $PD$ from the embeddings. We cast this as a Linear Assignment Problem (LAP; Burkard & Cela (1999)) solved by the Hungarian algorithm (Kuhn, 1955), which is invariant to any nonzero scalar factor. Specifically, we first restrict to the shared vocabulary and form the matrix of absolute cosine similarities between the columns of $W_{A,\text{emb}}$ and $W_{B,\text{emb}}$. Then, LAP is applied to obtain the permutation $P$. Next, we use the signs of the cosine similarities at the matched pairs to reconstruct the signature matrix $D$. This approach effectively reconstruct corresponding transformations due to its low demand of additional parameters in the detection process.
Robust Recovery of Weight Similarities
Despite accounting for permutation and signature manipulations, the orthogonal transformations $U_{B,i}^{(l)}$ remain challenging. They introduce a substantial nuisance parameter burden: if $W_{A,i}^{(l)}\in\mathbb{R}^{d\times n}$, then orthogonal $U_{B,i}^{(l)}\in\mathbb{R}^{d\times d}$ contributes $d^{2}$ parameters, which is prohibitive when $d\approx n$ and can undermine robustness by adding parameters to detections (Simmons et al., 2011). We therefore use Central Kernel Alignment (CKA) as a parameter-free similarity metric: by Theorem 3.1, CKA is invariant to orthogonal transforms and constant rescaling. While this invariance does not extend to semi-orthogonal $U_{B,i}^{(l)}$ (e.g., induced by pruning), CKA remains effective to a meaningful extent in such cases, relieving the need to explicitly reconstruct $U_{B,i}^{(l)}$. To further mitigate biases (Gretton et al., 2005; Murphy et al., 2024), we adopt the unbiased variant (Song et al., 2007), termed UCKA (see Appendix A.2).
6 Experiments
In this section, we present a series of experiments designed to rigorously evaluate our proposed model fingerprinting method. We begin in Section 6.1 by verifying its fundamental ability to distinguish between derived (offspring) and independent models. Next, in Section 6.2, we assess the critical risk of false positives by comparing its performance on 90 pairs of independent models against two state-of-the-art baselines, HuRef and REEF. In Section 6.3, we test the method’s robustness against a wide array of common post-training modifications using a comprehensive suite of 60 offspring models. Finally, in Section 6.4, we provide an overall performance comparison, leveraging ROC curves and other metrics to demonstrate our method’s superior discriminative power.
Baselines. We compare our method against two advanced baselines: the weight-based method HuRef (Zeng et al., 2024), and the representation-based method REEF (Zhang et al., 2025).
6.1 Effectiveness Verification
We first established our method’s core effectiveness in identifying model lineage. We collected eight offspring models for both LLaMA2-7B and LLaMA2-13B, alongside eight independent models for each size. We then calculated the similarity between the base model and these two groups.
As shown in Table 1, the results demonstrate a clear distinction. Offspring models exhibited exceptionally high similarity to their respective base models, whereas the similarity scores between independent models were negligible, forming a strong basis for intellectual property protection.
Table 1: Similarity (%) of various LLMs to LLaMA2-7B and LLaMA2-13B base models. Offspring models consistently show high similarity scores (light red), while independent models have negligible similarity (light green), demonstrating the method’s discriminative power.
Base Model: LLaMA2-7B Base Model: LLaMA2-13B
Offspring Sim Independent Sim Offspring Sim Independent Sim
WizardMath-7b 99.99 Baichuan-7b 0.58 Selfrag_llama2_13b 99.99 Baichuan-13b 0.17
Selfrag_7b 99.98 Mistral-7b 0.63 Nous-Hermes-13b 99.98 Baichuan2-13b 0.23
Vicuna-7b 99.95 OLMo-7b 0.46 Llama2-13b-orca 99.98 OLMo2-13b 0.21
Llama2-7b-Chat 99.93 Qwen-7b 0.19 Vicuna-13b 99.93 Qwen-14b 0.20
Finance-7b 99.93 InternLM-7b 0.43 Llama2-13b-Chat 99.92 Qwen3-14b 0.41
Llama2-7b-32K 99.88 MPT-7b 0.04 Firefly-llama2-13b 99.81 Jais-13b 0.03
Guanaco-7b 99.80 LLaMA-7b 0.81 Llama2-koen-13b 97.76 LLaMA-13b 0.74
Llama2-ko-7b 96.79 OpenLlama-7b 0.71 Llama2-13b-Estopia 96.60 OpenLlama-13b 0.51
6.2 False Positive Risk Evaluation on 90 Unrelated Pairs
A critical requirement for any reliable fingerprinting method is an extremely low false positive rate. To evaluate this risk, we collected 10 independent 7B LLMs and 10 independent 13B LLMs, forming 45 unique pairs for each size. We then computed the pairwise similarity (%) for all 90 pairs using our method and compared the results with those from REEF and HuRef.
The heatmaps in Figure 2 illustrate our method’s superior performance in avoiding false positives. For independent pairs, our method yielded mean similarity scores of just 0.49 (7B) and 0.26 (13B). These scores are nearly an order of magnitude lower than HuRef (means of 3.56 and 2.17) and two orders of magnitude lower than REEF (means of 42.47 and 47.44).
Notably, REEF frequently produces dangerously high similarity scores for unrelated models, with many pairs exceeding 80 and some even surpassing 95. Such high values could easily lead to false accusations of model theft. In contrast, the maximum similarity scores for our method were merely 1.5 (7B) and 0.8 (13B), reaffirming its significantly lower risk of false positives.
Figure 2: Pairwise similarity (%) heatmaps for independent 7B (top row) and 13B (bottom row) LLMs. The columns compare our proposed method against two baselines, REEF and HuRef. Our method consistently yields near-zero similarity scores for all independent model pairs, indicating a significantly lower risk of false positives. In contrast, REEF often produces high similarity scores (>80>80), which could easily lead to false accusations of model theft.
6.3 Robustness Verification on 60 Offspring Models
Table 2: Absolute Z-scores ($\uparrow$) of positive samples for HuRef, REEF, and our method. The full names of abbreviations are provided in Appendix C.
SFT
Base Llama2-7B Llama2-13B
Offspring Vicuna Finance Selfrag 32K Wizard Guanaco Vicuna Hermes Estopia Firefly
HuRef 44.22 44.20 44.47 44.22 44.51 44.12 44.25 44.48 43.09 44.51
REEF 1.95 1.94 1.95 1.88 1.95 1.94 1.94 1.95 1.91 1.95
Ours 355.09 355.02 355.20 354.84 355.23 354.55 355.02 355.20 343.14 354.59
Continue Pretrain
Base Llama2-7B
Gemma
-2B
Gemma
-7B
Qwen2.5
-7B
Qwen2
-7B
Llama2-70B
Offspring
Llemma
Code
Python
Code
Code
Math
Coder
Math
Code
Python
HuRef 8.15 9.41 9.19 28.97 39.51 0.28 0.57 1.01 3.43 2.78
REEF 1.66 1.31 1.77 0.32 1.52 1.04 1.29 1.09 1.93 1.92
Ours 250.32 250.78 253.28 198.18 268.72 177.57 135.17 183.49 241.12 232.92
Upcycling
Base
Mistral
-7B
Llama3
-8B
Llama2-7B
Qwen
-1.8B
Minicpm
-2B
Offspring Mixtral
MoE
v2
MoE4
MoE
3B
MoE2
MoE3B
-SFT
MoE2
-SFT
MoE4
-SFT
Qwen1.5
MoE
Minicpm
MoE
HuRef 7.97 42.50 24.58 21.59 24.91 21.59 24.91 24.58 4.03 11.39
REEF 1.47 0.47 0.64 0.64 0.62 0.63 0.63 0.62 0.57 1.49
Ours 239.01 332.12 332.23 326.49 332.12 326.49 332.12 332.23 173.36 150.15
Multi Modal
Base Llama2-7B
Qwen2
-7B
Qwen-7B
Qwen2.5
-7B
Qwen2.5
-3B
Llama3
-8B
Llama2
-13B
Offspring LLaVA Video VL Audio Audio2 VL VL VL Next LLaVA
HuRef 44.06 44.03 39.94 38.79 21.59 37.41 24.28 23.95 44.30 44.09
REEF 0.40 0.36 0.84 0.19 0.45 0.77 0.48 0.61 0.26 0.07
Ours 354.98 354.98 342.72 339.86 317.97 336.55 290.08 298.46 355.05 354.91
RL
Base
Open-
llama3B
Qwen2.5
-7B
Qwen2.5
-1.5B
Mixtral Mistral-7B
Minicpm
-2B
Qwen3
-4B
Chatglm
-6B
Llama3
-8B
Offspring RLHF Reason Zero DPO DPO Dolphin DPO GRPO RLHF DPO
HuRef 44.52 44.58 44.58 44.57 44.53 44.54 44.58 44.58 44.58 44.52
REEF 1.94 1.93 1.94 1.75 1.48 1.21 1.92 1.78 1.96 1.96
Ours 355.23 355.27 355.27 355.23 355.23 355.23 355.27 355.27 355.27 355.23
Pruning
Base Llama-3-8B Llama2-7B
Offspring
Minitron
-Depth
Minitron
-Width
Llama3
-1B
Llama3
-3B
Sheared
2.7B-P
Sheared
2.7B-S
Sheared
2.7B
Sheared
1.3B-P
Sheared
1.3B
Sheared
1.3B-S
HuRef 28.29 22.23 0.33 0.73 22.88 16.00 15.86 10.06 7.64 7.64
REEF 0.52 0.53 1.15 0.92 1.75 1.78 1.77 1.80 1.79 1.79
Ours 344.07 343.00 12.14 106.29 328.81 312.44 312.80 317.79 297.50 297.50
Base LLMs often undergo substantial post-training modifications, including SFT, continued pretraining, reinforcement learning (RL), pruning, upcycling, and multi-modal adaptation. These processes can significantly alter model parameters, making robustness a critical attribute for any fingerprinting technique. To rigorously evaluate our method’s robustness, we curated a diverse test suite of 60 positive pairs, each consisting of a base model and a derived offspring model. This suite includes 10 pairs for each of the six modification categories listed above.
The effectiveness of a fingerprinting method hinges on its ability to distinguish between related (positive) and unrelated (negative) pairs. However, absolute similarity scores (%) can be misleading. A method yielding scores of 10 for a positive pair and 0.1 for negative pairs is far more discriminative than one producing scores of 90 and 40. To accurately measure the discriminative power, we must quantify how statistically different a positive pair’s similarity is from the distribution of negative pairs. For this, we use the absolute Z-score (|Z||Z|) [^3] In this context, the absolute Z-score is mathematically equivalent to the Mahalanobis distance.. This metric measures how many standard deviations a positive pair’s similarity is from the mean of the negative pairs’ similarities. A larger $|Z|$ indicates that the positive pair is more likely to be a statistical outlier relative to the population of unrelated pairs and thus more highly separable.
The results, presented in Table 2, highlight the superior robustness of our method. HuRef loses effectiveness under heavy modifications like extensive continued pretraining (e.g., its $|Z|$ drops to 0.57 for Qwen2.5-coder). REEF consistently yields low $|Z|$, often below 2.0, indicating poor separability. In stark contrast, our method maintains remarkably high $|Z|$ across all 60 positive pairs, demonstrating its resilience against a wide array of demanding post-training modifications.
6.4 Overall Evaluation
Table 3: Detailed performance comparison of fingerprinting methods across various post-training techniques. Our method consistently achieves perfect scores (1.0) on all classification metrics (AUC, pAUC, TPR@1%FPR) and maintains a significantly larger separation margin ($\overline{|Z|}$) across all scenarios. CPT: Continual Pre-Training, UP: Upcycling, $M$: Multi-modal, PR: Pruning.
Method Metric SFT CPT UP $M$ RL PR All
HuRef $\overline{|Z|}$ $\uparrow$ 43.748 10.331 20.805 36.244 44.559 13.166 28.142
AUC $\uparrow$ 1.000 0.879 0.999 1.000 1.000 0.866 0.957
pAUC $\uparrow$ 1.000 0.656 0.978 1.000 1.000 0.800 0.906
TPR@1%FPR $\uparrow$ 1.000 0.500 0.900 1.000 1.000 0.800 0.867
REEF $\overline{|Z|}$ $\uparrow$ 1.936 1.384 0.777 0.443 1.788 1.381 1.285
AUC $\uparrow$ 1.000 0.857 0.508 0.648 0.963 0.692 0.778
pAUC $\uparrow$ 1.000 0.211 0.000 0.000 0.658 0.300 0.362
TPR@1%FPR $\uparrow$ 1.000 0.200 0.000 0.000 0.600 0.000 0.300
Ours $\overline{|Z|}$ $\uparrow$ 353.788 219.155 287.634 334.556 355.250 267.233 302.936
AUC $\uparrow$ 1.000 1.000 1.000 1.000 1.000 1.000 1.000
pAUC $\uparrow$ 1.000 1.000 1.000 1.000 1.000 1.000 1.000
TPR@1%FPR $\uparrow$ 1.000 1.000 1.000 1.000 1.000 1.000 1.000
To synthesize our findings, we conducted a comprehensive evaluation against HuRef and REEF using the full dataset of 60 positive and 90 negative model pairs.
As illustrated in Figure 1, our method demonstrates vastly superior performance. The Receiver Operating Characteristic (ROC) curve (left) shows that our method achieves a perfect Area Under the Curve (AUC) of 1.0, indicating flawless discrimination. This starkly outperforms both HuRef ($AUC=0.957$) and REEF ($AUC=0.778$).
In practical applications, preventing false accusations of model theft is paramount, making performance at a very low False Positive Rate (FPR) essential. We therefore employ two stricter metrics: the partial AUC for FPR <0.05<0.05 (pAUC) and the True Positive Rate at a 1% FPR (TPR@1%FPR).
Table 3 details the performance breakdown across post-training techniques. Our method consistently achieves perfect scores (1.0) on all classification metrics (AUC, pAUC, and TPR@1%FPR) across all categories. In contrast, the baseline methods show significant limitations. REEF’s performance collapses in several scenarios, with pAUC and TPR@1%FPR scores falling to 0.0 for Upcycling and Multi-modal. HuRef also shows vulnerability, with its TPR@1%FPR dropping to 0.500 under Continual Pre-Training. Our method’s perfect classification scores, combined with its substantially larger average absolute Z-score ($\overline{|Z|}$), underscore its superior robustness and reliability.
7 Conclusions
In this paper, we propose a training-free fingerprinting method for LLM identification. Our approach does not impair LLM’s general capability while exhibiting robustness against fine-tuning, extensive continued pretraining, reinforcement learning, multimodal extension, pruning, and upcycling, and simultaneously avoids the risk of false positives. Experiments on a testbed comprising 150 pairs of LLMs demonstrate the effectiveness of our method.
Acknowledgement
This work is sponsored by the National Key Research and Development Program of China (No. 2023ZD0121402) and the National Natural Science Foundation of China (NSFC) grant (No.62576211).
Appendix A Definitions and Notations
A.1 LLM Architecture
Weights
Here we provide notations for the rest of model A’s weights. In addition to the notations in Section 3, we further denote s, $W_{A,\text{lm}}$ as the language model head, and $W_{A,\text{V}}^{(l)},W_{A,\text{O}}^{(l)},\mathbb{W}_{A,\text{ffn}}^{(l)}$ the V,O matrices and FFN at the $l$-th layer, and
$\mathbb{W}_{A,\text{others}}$ $=\mathbb{W}_{A}/\mathbb{W}_{A,\text{partial}}$
$=\{W_{A,\text{lm}}\}\bigcup_{l=1}^{L}\mathbb{W}_{A,\text{ffn}}^{(l)}\bigcup_{l=1}^{L}\{W_{A,i}^{(l)}\mid i\in\{\text{V},\text{O}\}\}$
as the rest of model A’s weights.
Functions
Here we provide definitions for functions in an LLM.
Definition A.1 (RMSNorm).
Given the input matrix $X\in\mathbb{R}^{m\times n}$ of model A, the learnable parameters $\omega_{A}^{(l)}=(\omega_{A,1}^{(l)},\dots,\omega_{A,n}^{(l)})\in\mathbb{R}^{n}$ at the $l$-th layer, and a constant $\epsilon_{A}\in\mathbb{R}$ (this constant is usually shared across all layers of an LLM), RMSNorm at layer $l$ is a function that
$\text{RMSNorm}(X,\omega_{A}^{(l)},\epsilon_{A})=\text{diag}^{-\frac{1}{2}}(XX^{\top}\mathbf{1}_{m}+\epsilon_{A}\mathbf{1}_{m})\cdot X\cdot\text{diag}(\omega_{A}^{(l)}).$ (6)
Definition A.2 (Self-Attention with RoPE).
Given an input $X$ to model A, the self-attention mechanism at layer $l$ is a function $f_{\text{attn}}$ that
$f_{\text{attn}}^{(l)}(X)=\text{AttnScore}(X,W_{A,\text{Q}}^{(l)},W_{A,\text{K}}^{(l)},\theta)\cdot XW_{A,\text{V}}^{(l)\top}\cdot W_{A,\text{O}}^{(l)\top}$ (7)
where AttnScore is the attention score function with RoPE, i.e.
$\text{AttnScore}(X,W_{A,\text{Q}}^{(l)},W_{A,\text{K}}^{(l)},\theta)=\text{softmax}\Big(\frac{1}{\sqrt{d}}\text{RoPE}(XW_{A,\text{Q}}^{(l)\top},\theta)\cdot\text{RoPE}(XW_{A,\text{K}}^{(l)\top},\theta)^{\top}\Big),$ (8a)
$\text{RoPE}(XW_{A,\text{Q}}^{(l)\top},\theta)=[x_{1}^{\top}W_{A,\text{Q}}^{(l)\top}R_{\theta}(0),\dots,x_{m}^{\top}W_{A,\text{Q}}^{(l)\top}R_{\theta}(m-1)],$ (8b)
$\text{RoPE}(XW_{A,\text{K}}^{(l)\top},\theta)=[x_{1}^{\top}W_{A,\text{K}}^{(l)\top}R_{\theta}(0),\dots,x_{m}^{\top}W_{A,\text{K}}^{(l)\top}R_{\theta}(m-1)].$ (8c)
Here $x_{i}^{\top}\in\mathbb{R}^{n}$ is the i-th row of $X$, $\theta$ is the frequency base for RoPE and $R_{\theta}$ is the rotation matrix (Su et al., 2024).
RoPE’s definition in Section A.1 is slightly different from their implementations, but still an equivalent version to Su et al. (2024).
Definition A.3 (Feed-forward Networks).
Given the input $X$, the feed-forward network which adopts the SwiGLU MLP is the function $f^{(l)}_{\text{ffn}}$ defined by
$f^{(l)}_{\text{ffn}}(X)\;=\;\Big(\text{SiLU}\big(X\,{W^{(l)\top}_{A,\text{gate}}}\big)\odot\big(X\,{W^{(l)\top}_{A,\text{up}}}\big)\Big)\;{W^{(l)\top}_{A,\text{down}}}$,
where $W^{(l)}_{A,\text{up}}$, $W^{(l)}_{A,\text{gate}}$, $W^{(l)}_{A,\text{down}}$ are bias-free weight matrices, $\odot$ denotes the Hadamard product, and $\text{SiLU}(z)=z\,\sigma(z)$ with $\sigma(z)=\frac{1}{1+e^{-z}}$. We write $\mathbb{W}^{(l)}_{A,\text{ffn}}=\big\{W^{(l)}_{A,\text{up}},\,W^{(l)}_{A,\text{gate}},\,W^{(l)}_{A,\text{down}}\big\}$.
We abuse the notations in Section A.1 by denoting $\text{RMSNorm}_{\text{attn}}^{(l)}(X)\triangleq\text{RMSNorm}(X,\omega_{A,\text{attn}}^{(l)},\epsilon_{A})$ as the attention norm and $\text{RMSNorm}_{\text{ffn}}^{(l)}(X)\triangleq\text{RMSNorm}(X,\omega_{A,\text{ffn}}^{(l)},\epsilon_{A})$ the ffn norm at layer $l$ of model A. Then, the definition of a Transformer block at layer $l$ of model A can be summartized as follows.
Definition A.4 (Transformer Block).
Combining all the components with residual connections, the transformer block at the $l$-th layer of model A is a function $f_{\text{Transformer}}^{(l)}$ that
$f_{\text{Transformer}}^{(l)}=\Big(\text{I}+f_{\text{ffn}}^{(l)}\circ\text{RMSNorm}^{(l)}_{\text{ffn}}\Big)\circ\Big(\text{I}+f_{\text{attn}}^{(l)}\circ\text{RMSNorm}^{(l)}_{\text{attn}}\Big)$. (9)
where I is the identity mapping, $f_{\text{attn}}^{(l)}$ and $f_{\text{ffn}}^{(l)}$ are the attention and FFN at $l$-th layer respectively, $\text{RMSNorm}^{(l)}_{\text{ffn}}$ and $\text{RMSNorm}^{(l)}_{\text{attn}}$ are RMSNorm functions for these two components.
Definition A.5 (LLM Architecture).
Given an word embedding $X$, the last hidden state $Y$ of an $L$-layered LLM A is
$Y=f_{\text{Transformer}}^{(L)}\circ f_{\text{Transformer}}^{(L-1)}\circ\dots\circ f_{\text{Transformer}}^{(L-1)}(X)\cdot W_{A,\text{lm}}^{\top}.$ (10)
A.2 Central Kernel Alignment (CKA)
Given two input matrices $X_{1}\in\mathbb{R}^{m\times n_{1}}$ and $X_{2}\in\mathbb{R}^{m\times n_{2}}$, CKA is a function mapping paired input matrices to $[0,1]$, and
$\text{CKA}(X_{1},X_{2})=\frac{\text{HSIC}(K_{X_{1}},K_{X_{2}})}{\sqrt{\text{HSIC}(K_{X_{1}},K_{X_{1}})\cdot\text{HSIC}(K_{X_{2}},K_{X_{2}})}}$ (11)
where $\text{HSIC}(K_{X_{1}},K_{X_{2}})=\frac{1}{(m-1)^{2}}\text{tr}(K_{X_{1}}H_{m}K_{X_{2}}H_{m})$, $(K_{X_{1}})_{ij}=k((X_{1})_{i},(X_{1})_{j})$, $(K_{X_{2}})_{ij}=k((X_{2})_{i},(X_{2})_{j})$ are kernel matrices with the kernel function $k(\cdot,\cdot):\mathbb{R}^{m}\times\mathbb{R}^{m}\rightarrow\mathbb{R}_{+}$, and $H_{m}=I_{m}-\frac{1}{m}\mathrm{1_{m}}\mathrm{1_{m}}^{\top}$ is the centering matrix. The linear kernels give $K_{X}=XX^{\top}$ for a matrix $X$. To reduce the finite-sample bias of Eq. (6), we use the unbiased HSIC estimator (Song et al., 2007). Let $\tilde{K}_{X_{i}}$ be $K_{X_{i}}$ with its diagonal set to zero, i.e., $(\tilde{K}_{X_{i}})_{ii}=0$. With $\mathrm{1}_{m}$ the all-ones vector of length $m$, the unbiased HSIC is
$\text{HSIC}_{u}(\tilde{K}_{X_{1}},\tilde{K}_{X_{2}})=\frac{1}{m(m-3)}\!\left[\mathrm{tr}(\tilde{K}_{X_{1}}\tilde{K}_{X_{2}})+\frac{(\mathrm{1}_{m}^{\top}\tilde{K}_{X_{1}}\mathrm{1}_{m})(\mathrm{1}_{m}^{\top}\tilde{K}_{X_{2}}\mathrm{1}_{m})}{(m-1)(m-2)}-\frac{2}{m-2}\,\mathrm{1}_{m}^{\top}\tilde{K}_{X_{1}}\tilde{K}_{X_{2}}\mathrm{1}_{m}\right].$
We then define
$\mathrm{UCKA}(X_{1},X_{2})=\frac{\text{HSIC}_{u}(\tilde{K}_{X_{1}},\tilde{K}_{X_{2}})}{\sqrt{\text{HSIC}_{u}(\tilde{K}_{X_{1}},\tilde{K}_{X_{1}})\cdot\text{HSIC}_{u}(\tilde{K}_{X_{2}},\tilde{K}_{X_{2}})}},$ (13)
which preserves Theorem 3.1 and yields values in [1,1][-1,1].
Appendix B Proofs and Discussions
B.1 Proof for Theorem 3.1
We give the proof based on the definition of CKA with linear kernels in Appendix A.2. Given input matrices $X_{1},X_{2}$, orthogonal matrices $U_{1},U_{2}$ and constants $c_{1},c_{2}$, the kernel functions yield
$K_{c_{1}X_{1}U_{1}}=c_{1}X_{1}U_{1}U_{1}^{\top}X_{1}^{\top}c_{1}=c_{1}^{2}X_{1}X_{1}^{\top}=c_{1}^{2}K_{X_{1}},$
$K_{c_{2}X_{2}U_{2}}=c_{2}X_{2}U_{2}U_{2}^{\top}X_{2}^{\top}c_{2}=c_{2}^{2}X_{2}X_{2}^{\top}=c_{2}^{2}K_{X_{2}}.$
Therefore, corresponding HSIC results are
$\text{HSIC}(K_{c_{1}X_{1}U_{1}},K_{c_{2}X_{2}U_{2}})=\frac{c_{1}^{2}c_{2}^{2}}{(m-1)^{2}}\text{tr}(K_{X_{1}}H_{m}K_{X_{2}}H_{m})=c_{1}^{2}c_{2}^{2}\text{HSIC}(K_{X_{1}},K_{X_{2}}),$
$\text{HSIC}(K_{c_{1}X_{1}U_{1}},K_{c_{1}X_{1}U_{1}})=\frac{c_{1}^{4}}{(m-1)^{2}}\text{tr}(K_{X_{1}}H_{m}K_{X_{1}}H_{m})=c_{1}^{4}\text{HSIC}(K_{X_{1}},K_{X_{1}}),$
$\text{HSIC}(K_{c_{2}X_{2}U_{2}},K_{c_{2}X_{2}U_{2}})=\frac{c_{2}^{4}}{(m-1)^{2}}\text{tr}(K_{X_{2}}H_{m}K_{X_{2}}H_{m})=c_{2}^{4}\text{HSIC}(K_{X_{2}},K_{X_{2}}).$
Hence, it follows that
$\text{CKA}(c_{1}X_{1}U_{1},c_{2}X_{2}U_{2})$ $=\frac{\text{HSIC}(K_{c_{1}X_{1}U_{1}},K_{c_{2}X_{2}U_{2}})}{\sqrt{\text{HSIC}(K_{c_{1}X_{1}U_{1}},K_{c_{1}X_{1}U_{1}})\cdot\text{HSIC}(K_{c_{2}X_{2}U_{2}},K_{c_{2}X_{2}U_{2}})}}$
$=\frac{c_{1}^{2}c_{2}^{2}\text{HSIC}(K_{X_{1}},K_{X_{2}})}{c_{1}^{2}c_{2}^{2}\sqrt{\text{HSIC}(K_{X_{1}},K_{X_{1}})\cdot\text{HSIC}(K_{X_{2}},K_{X_{2}})}}$
$=\frac{\text{HSIC}(K_{X_{1}},K_{X_{2}})}{\sqrt{\text{HSIC}(K_{X_{1}},K_{X_{1}})\cdot\text{HSIC}(K_{X_{2}},K_{X_{2}})}}$
$=\text{CKA}(X_{1},X_{2}).$
B.2 Proof for Section 4.2
The goal is to show the manipulation $T$ commutes with $f$, i.e. $f\circ T=T\circ f$. If $f=f_{n}\circ\dots\circ f_{1}$, then it suffices to show
$f_{n}\circ\dots\circ f_{1}\circ T=T\circ f_{n}\circ\dots f_{1}.$
Since
$f_{1}\circ T=T\circ f_{1},f_{2}\circ T=T\circ f_{2},\dots,f_{n}\circ T=T\circ f_{n},$
It is direct that
$f\circ T=T\circ f.$
B.3 Proof for Theorem 4.3 and Discussions on Other Input Manipulations
We first show a proof for the claims in Theorem 4.3. The core lies in preserving the diagonal structure of RMSNorm parameters in Section A.1. Given the input $X$ to layer $l$, the manipulations change it to $cXPD$ and fed the manipulated input into the attention norm of model B. In order to recover the manipulations on inputs after RMSNorm, i.e. for any manipulation $T$ in the input $X$ (an example is $T(X)=cXPD$),
$\text{RMSNorm}_{\text{attn}}^{(l)}\circ T=T\circ\text{RMSNorm}_{\text{attn}}^{(l)},$ (14)
it suffices to show
$\exists\omega_{B,\text{attn}}^{(l)}\text{ and }\epsilon_{B}\text{, s.t. }\text{RMSNorm}(X,\omega_{A,\text{attn}}^{(l)},\epsilon_{A})\cdot cPD=\text{RMSNorm}(cXPD,\omega_{B,\text{attn}}^{(l)},\epsilon_{B})$
since the learnable parameters and norm epsilon can be modified to satisfy Equation 14 and Section 4.2. By setting $\epsilon_{B}=c^{2}\epsilon_{A}$ and $\text{diag}(\omega_{B,\text{attn}}^{(l)})=cD^{\top}P^{\top}\text{diag}(\omega_{A,\text{attn}}^{(l)})PD$, one has
$\text{RMSNorm}(cXPD,\omega_{B,\text{attn}}^{(l)},\epsilon_{B})$
$=\text{diag}^{-\frac{1}{2}}\left((cXPD)(cXPD)^{\top}\mathbf{1}_{m}+\epsilon_{B}\mathbf{1}_{m}\right)\cdot cXPD\cdot\text{diag}(\omega_{B,\text{attn}}^{(l)})$
$=\text{diag}^{-\frac{1}{2}}(c^{2}XX^{\top}\mathbf{1}_{m}+c^{2}\epsilon_{A}\mathbf{1}_{m})\cdot cXPD\cdot cD^{\top}P^{\top}\text{diag}(\omega_{A,\text{attn}}^{(l)})PD$
$=c\cdot\text{diag}^{-\frac{1}{2}}(XX^{\top}\mathbf{1}_{m}+\epsilon_{A}\mathbf{1}_{m})\cdot X\cdot\text{diag}(\omega_{A,\text{attn}}^{(l)})\cdot PD$
$=\text{RMSNorm}(X,\omega_{A,\text{attn}}^{(l)})\cdot cPD.$
The proof still holds for partial permutations $P$, i.e. $P$ is rectangular with number of rows more than number of columns, since $cD^{\top}P^{\top}\text{diag}(\omega_{A,\text{attn}}^{(l)})PD$ is still a diagonal matrix. Therefore, $R_{B,\text{emb}}=cPD$ under Section 4.2. However, for more general manipulations, Section 4.2 does not hold for RMSNorm. We give two examples for illustration.
First, orthogonal transformations on the input generally fail to pass through RMSNorm. Given any orthogonal matrices $U$ as the manipulation on $X$, i.e., $T(X)=XU$, if one attempts to keep Equation 14, then it requires
$\exists\omega^{(l)}\text{ and }\epsilon,\text{ s.t. }\text{RMSNorm}(XU,\omega^{(l)},\epsilon)=\text{RMSNorm}(X,\omega_{A,\text{attn}}^{(l)},\epsilon_{A})U.$
However,
$\text{RMSNorm}(XU,\omega,\epsilon)=\text{diag}^{-\frac{1}{2}}(XX^{\top}\mathbf{1}_{m}+\epsilon\mathbf{1}_{m})\cdot XU\cdot\text{diag}(\omega^{(l)}).$
Hence, it is required that
$U\cdot\text{diag}(\omega^{(l)})=\text{diag}(\omega_{A,\text{attn}}^{(l)})U$
which suggests $U^{\top}\text{diag}(\omega_{A,\text{attn}}^{(l)})U$ is a diagonal matrix. Nevertheless, this property generally does not hold if $U$ is not a (partial) permutation matrix $P$ or a signature matrix $D$.
Second, non-orthogonal transformations on inputs generally fail to pass through RMSNorm. Given an arbitrary (invertible) transformation $M$, Section 4.2 and Equation 14 require that
$\exists\omega^{(l)}\text{ and }\epsilon,\text{ s.t. }\text{RMSNorm}(XM,\omega^{(l)},\epsilon)=\text{RMSNorm}(X,\omega_{A,\text{attn}}^{(l)},\epsilon_{A})M.$
However, since
$\text{RMSNorm}(XM,\omega^{(l)},\epsilon)=\text{diag}^{-\frac{1}{2}}(XMM^{\top}X^{\top}\mathbf{1}_{m}+\epsilon\mathbf{1}_{m})\cdot XM\cdot\text{diag}(\omega^{(l)}),$
any $M$ that do not satisfy $MM^{\top}=c^{\prime}I$ where $c^{\prime}$ is a constant and $I$ is the identity matrix can hardly recover the manipulations due to the nonlinearity in norm functions. Moreover, similar to the case with orthogonal transformations, $M\cdot\text{diag}(\omega^{(l)})=\text{diag}(\omega_{A,\text{attn}}^{(l)})M$ also requires $M$ to be (partial) permutation or signature matrices (or a combination of the both).
Third, we additionally clarify that although the combination of (partial) permutation and signature matrices can be in any order, it is reasonable to fix it as $R_{B,\text{emb}}=cPD$. To be specific, we clarify that for any (partial) permutation matrix $P^{\prime}$ and signature matrix $D^{\prime}$, there exists a (partial) permutation matrix $P$ and a signature $D$ such that
$D^{\prime}P^{\prime}=PD$
The proof is straightforward. Define $P$ entry-wise by
$P_{ij}:=\bigl|(D^{\prime}P^{\prime})_{ij}\bigr|\in\{0,1\}.$
Since left-multiplication by $D^{\prime}$ only flips signs, each column of $D^{\prime}P^{\prime}$ has at most one nonzero entry. Hence, $P$ is a partial permutation matrix. For each column $j$, if that column of $D^{\prime}P^{\prime}$ has its unique nonzero at row $i$, set
$D_{jj}:=\operatorname{sgn}\bigl((D^{\prime}P^{\prime})_{ij}\bigr)\in\{\pm 1\}.$
If the column is entirely zero, choose $D_{jj}\in\{\pm 1\}$ arbitrarily. Then for all $i,j$,
$(PD)_{ij}=P_{ij}D_{jj}=\bigl|(D^{\prime}P^{\prime})_{ij}\bigr|\cdot\operatorname{sgn}\bigl((D^{\prime}P^{\prime})_{ij}\bigr)=(D^{\prime}P^{\prime})_{ij},$
where we interpret $\operatorname{sgn}(0)=1$. Therefore, $PD=D^{\prime}P^{\prime}$.
Last, we clarify that the conclusions in Theorem 4.3 can generalize to the case where model A and model B may not share the same architecture. It is a direct result from the fact that the two LLMs have shared tokens and that the pruning over dimensions of a model’s word embeddings can be viewed as a partial permutation matrix.
B.4 Proof for Theorem 4.4 and Related Discussions
We first provide the proof for Theorem 4.4 based on Section A.1.
We denote by $X$ the original input to self-attention, and $X^{\prime}$ the manipulated one. Then, Theorem 4.3 suggests that
$X^{\prime}=cXPD$
Therefore, the attention score of model B becomes
$\text{AttnScore}(cXPD,W_{B,\text{Q}}^{(l)},W_{B,\text{K}}^{(l)},\theta)=\text{softmax}\Big(\frac{1}{\sqrt{d}}\text{RoPE}(cXPDW_{B,\text{Q}}^{(l)\top},\theta)\cdot\text{RoPE}(cXPDW_{B,\text{K}}^{(l)\top},\theta)^{\top}\Big).$
To keep attention scores of model B identical to that of model A, it is required that
$\text{RoPE}(cXPDW_{B,\text{Q}}^{(l)\top},\theta)\cdot\text{RoPE}(cXPDW_{B,\text{K}}^{(l)\top},\theta)^{\top}=\text{RoPE}(XW_{A,\text{Q}}^{(l)\top},\theta)\cdot\text{RoPE}(XW_{A,\text{K}}^{(l)\top})^{\top}.$
By Equation 8b and Equation 8c, this translates into
$c^{2}x_{i}^{\top}PDW_{B,\text{Q}}^{(l)\top}R_{\theta}(i)R_{\theta}^{\top}(j)W_{B,\text{K}}^{(l)}D^{\top}P^{\top}x_{j}=x_{i}^{\top}W_{A,\text{Q}}^{(l)\top}R_{\theta}(i)R_{\theta}^{\top}(j)W_{A,K}^{(l)}x_{j}$
Therefore, the recovery of model A’s attention score requires column-wise transformations to eliminate the input-related manipulations passed by RMSNorm, i.e.
$W_{B,\text{Q}}^{(l)}=c^{-1}W_{A,\text{Q}}^{(l)}PD,\quad W_{B,\text{K}}^{(l)}=c^{-1}W_{A,\text{K}}^{(l)}PD$ (15)
which suggests
$R_{B,\text{Q}}^{(l)}=R_{B,\text{K}}^{(l)}=c^{-1}PD.$
Although Equation 15 recovers attention scores, the Q,K matrices can still be modified without changing attention scores. An example is, given a rotation matrix $R=R_{\theta}(k)$ with the same frequency base as RoPE, multiplying Q,K matrices with it does not change the attention scores, i.e.
$x_{i}^{\top}W_{B,\text{Q}}^{(l)\top}R_{\theta}(k)R_{\theta}(i)R_{\theta}^{\top}(j)R_{\theta}^{\top}(k)W_{B,\text{K}}^{(l)}x_{j}$ $=x_{i}^{\top}W_{B,\text{Q}}^{(l)\top}R_{\theta}(i+k)R_{\theta}^{\top}(j+k)W_{B,\text{K}}^{(l)\top}x_{j}$
$=x_{i}^{\top}W_{B,\text{Q}}^{(l)\top}R_{\theta}(i-j)W_{B,\text{K}}^{(l)}x_{j}$
$=x_{i}^{\top}W_{B,\text{Q}}^{(l)\top}R_{\theta}(i)R_{\theta}^{\top}(j)W_{B,\text{K}}^{(l)\top}x_{j}.$
Furthermore, this conclusion can be generalize to any rotation matrix with structures similar to RoPE rotation matrices. Given $R=\text{diag}(R(\psi_{0}),R(\psi_{1}),\dots)$ with $R(\psi_{k})=\begin{bmatrix}\cos\psi_{k}&-\sin\psi_{k}\\ \sin\psi_{k}&\cos\psi_{k}\end{bmatrix}$, multiplying Q,K matrices by $R$ keeps the attention scores since
$RR_{\theta}(i)R_{\theta}^{\top}(j)R^{\top}=R_{\theta}(i)R_{\theta}^{\top}(j).$ (16)
Hence, there are various rotation matrices to significantly change Q,K matrices but preserve attention scores. Since these rotation matrices are naturally orthogonal, we reformulate the row-wise transformations on Q,K matrices $L_{B,\text{Q}},L_{B,\text{K}}$ as $U_{B,\text{Q}}^{(l)}$ and $U_{B,\text{K}}^{(l)}$.
Therefore, we denote RoPE-related manipulations as
$W_{B,\text{Q}}^{(l)}=U_{A,\text{Q}}^{(l)}W_{B,\text{Q}}^{(l)},\quad W_{B,\text{K}}^{(l)}=U_{A,\text{K}}^{(l)}W_{B,\text{K}}^{(l)}.$
On the other hand, semi-orthogonal $U_{B,\text{Q}}^{(l)},U_{B,\text{K}}^{(l)}$ are also possible. This is because of the common practice in pruning may select different different rows of Q,K matrices, which suggests multiplying Q,K matrices with row-wise partial permutations.
Next, we show how Section 4.2 is satisfied at the self-attention mechanism based on Theorem 4.4. Let $W_{B,\text{V}}^{(l)}=W_{A,\text{V}}^{(l)}PD$ and $W_{B,\text{O}}^{(l)}=D^{\top}P^{\top}W_{A,\text{O}}^{(l)}$. Then,
$X^{\prime}W_{B,\text{V}}^{(l)\top}W_{B,\text{O}}^{(l)\top}$ $=cXPD\cdot D^{\top}P^{\top}W_{A,\text{V}}^{(l)\top}W_{A,\text{O}}^{(l)\top}PD$
$=XW_{A,\text{V}}^{(l)\top}W_{A,\text{O}}^{(l)\top}\cdot cPD$
recovers the manipulation over inputs in Equation 7.
B.5 How Potential Attacks in Section 4.5 Recovers the Output of Transformers
We provide an illustration based on Section A.1, Section A.1 and Section A.1. Previous parts have demonstrated how the manipulations on inputs pass through self-attention. Hence, it suffices to show how the manipulations pass through FFN and yield an output identical to model A after the language model head.
For layer $l$, let $W_{B,\text{gate}}^{(l)}=c^{-1}W_{A,\text{gate}}^{(l)}PD,W_{B,\text{gate}}^{(l)}=c^{-1}W_{A,\text{gate}}^{(l)}PD,W_{B,\text{down}}^{(l)}=cD^{\top}P^{\top}W_{A,\text{down}}^{(l)}$. We denote $X^{\prime}=cXPD$ as the manipulated input. Then for model B,
$f^{(l)}_{\text{ffn}}(X^{\prime})$
$=\Big(\text{SiLU}\big(X^{\prime}{W^{(l)\top}_{B,\text{gate}}}\big)\odot\big(X^{\prime}{W^{(l)\top}_{B,\text{up}}}\big)\Big){W^{(l)\top}_{B,\text{down}}}$
$=\Big(\text{SiLU}\big(cXPD\cdot D^{\top}P^{\top}W^{(l)\top}_{A,\text{gate}}\cdot c^{-1}\big)\odot\big(cXPD\cdot{D^{\top}P^{\top}W^{(l)\top}_{A,\text{up}}\cdot c^{-1}}\big)\Big){W^{(l)\top}_{B,\text{down}}}\cdot cPD$
$=\Big(\text{SiLU}\big(X{W^{(l)\top}_{A,\text{gate}}}\big)\odot\big(X{W^{(l)\top}_{A,\text{up}}}\big)\Big){W^{(l)\top}_{A,\text{down}}}\cdot cPD.$
By Section A.1 and Section A.1, the manipulation is propagated through transformer layers. At the language model head, we denote $W_{B,\text{lm}}=c^{-1}W_{A,\text{lm}}PD$ and abuse $X^{\prime}=cXPD$ as the manipulated input. Then,
$X^{\prime}W_{B,\text{lm}}^{(l)\top}$ $=cXPD\cdot D^{\top}P^{\top}W_{A,\text{lm}}^{(l)\top}\cdot c^{-1}$
$=XW_{A,\text{lm}}^{(l)\top}$
recovers the output of model A.
Appendix C Mapping of Model abbreviations
LABEL:tab:model_abbreviations details the mapping between the abbreviations used throughout the paper, their respective base models, and their full model names as cataloged on the Hugging Face Hub.
Table 4: Mapping of model abbreviations to their full Hugging Face model names.
Abbreviation Base Model Full Model Name
Vicuna Llama2-7B vicuna-7b-v1.5
Finance Llama2-7B llama-2-7b-finance
Selfrag Llama2-7B selfrag_llama2_7b
32K Llama2-7B LLaMA-2-7B-32K
Wizard Llama2-7B WizardMath-7B-V1.0
Guanaco Llama2-13B llama-2-7b-guanaco
Vicuna Llama2-13B vicuna-13b-v1.5
Hermes Llama2-13B Nous-Hermes-Llama2-13b
Estopia Llama2-13B LLaMA2-13B-Estopia
Firefly Llama2-13B firefly-llama2-13b
Llemma Llama2-7B llemma_7b
Code Llama2-7B CodeLlama-7b-hf
Python Llama2-7B CodeLlama-7b-Python-hf
Code Gemma-2B codegemma-2b
Code Gemma-7B codegemma-7b
Math Qwen2.5-7B Qwen2.5-Math-7B
Coder Qwen2.5-7B Qwen2.5-Coder-7B
Math Qwen2-7B Qwen2-Math-7B
Code Llama2-70B CodeLlama-70b-hf
Python Llama2-70B CodeLlama-70b-Python-hf
Mixtral Mistral-7B Nous-Hermes-2-Mixtral-8x7B-DPO
MoE v2 Llama3-8B LLaMA-MoE-v2-3_8B-2_8-sft
MoE4 Llama2-7B LLaMA-MoE-v1-3_5B-4_16
MoE 3B Llama2-7B LLaMA-MoE-v1-3_0B-2_16
MoE2 Llama2-7B LLaMA-MoE-v1-3_5B-2_8
MoE3B-SFT Llama2-7B LLaMA-MoE-v1-3_0B-2_16-sft
MoE2-SFT Llama2-7B LLaMA-MoE-v1-3_5B-2_8-sft
MoE4-SFT Llama2-7B LLaMA-MoE-v1-3_5B-4_16-sft
Qwen1.5 MoE Qwen-1.8B Qwen1.5-MoE-A2.7B
Minicpm MoE Minicpm-2B MiniCPM-MoE-8x2B
LLaVA Llama2-7B llava-v1.5-7b
Video Llama2-7B Video-LLaVA-7B-hf
VL Qwen2-7B Qwen2-VL-7B-Instruct
Audio Qwen-7B Qwen-Audio
Audio2 Qwen-7B Qwen2-Audio-7B
VL Qwen-7B Qwen-VL
VL Qwen2.5-7B Qwen2.5-VL-7B-Instruct
VL Qwen2.5-3B Qwen2.5-VL-3B-Instruct
Next Llama3-8B llama3-llava-next-8b-hf
LLaVA Llama2-13B llava-v1.5-13b
RLHF Open-llama3B hh_rlhf_rm_open_llama_3b
Reason Qwen2.5-7B Nemotron-Research-Reasoning-Qwen-1.5B
Zero Qwen2.5-1.5B Open-Reasoner-Zero-1.5B
DPO Mixtral Nous-Hermes-2-Mixtral-8x7B-DPO
DPO Mistral-7B Nous-Hermes-2-Mistral-7B-DPO
Dolphin Mistral-7B dolphin-2.6-mistral-7b-dpo
DPO Minicpm-2B MiniCPM-2B-dpo-bf16
GRPO Qwen3-4B Qwen3_Medical_GRPO
RLHF Chatglm-6B chatglm-fitness-RLHF
DPO Llama3-8B LLaMA3-iterative-DPO-final
Minitron-Depth Llama-3-8B Llama-3.1-Minitron-4B-Depth-Base
Minitron-Width Llama-3-8B Llama-3.1-Minitron-4B-Width-Base
Llama3-1B Llama3-8B Llama-3.2-1B
Llama3-3B Llama3-8B Llama-3.2-3B
Sheared 2.7B-P Llama2-7B Sheared-LLaMA-2.7B-Pruned
Sheared 2.7B-S Llama2-7B Sheared-LLaMA-2.7B-ShareGPT
Sheared 2.7B Llama2-7B Sheared-LLaMA-2.7B
Sheared 1.3B-P Llama2-7B Sheared-LLaMA-1.3B-Pruned
Sheared 1.3B Llama2-7B Sheared-LLaMA-1.3B
Sheared 1.3B-S Llama2-7B Sheared-LLaMA-1.3B-ShareGPT